# Effective Bertini theorems and zeros of $p$-adic forms of degree 7

### v1

@@ 2.1. A factorization algorithm
In this subsection, suppose $f(X,Y)\in k[X,Y]$ is monic in $X$ and has total degree $d$ (not necessarily homogeneous). Fix a positive integer $D<d$. The factorization algorithms found in [Kal95] and [CM06] underlying Theorem 2.1 request that $f(X,0)$ is squarefree. With slight modifications however, they still yield useful information when $f(X,0)$ is not squarefull, i.e. when considered over $\overline{k}$, $f(X,0)$ has at least one simple root. We recall the setup for the algorithm; see also [Kal95, §1], [CM06, §3.1].
@@
When $D=1$, a solution to (2.2) is equivalent to $f$ possessing a linear factor vanishing on $(\alpha_{0},0)$. This is essentially a modified version of [CM06, Lemma 3.1].
@@
Then $f$ factors over $\overline{k}$ as $f=g(X,Y)h(X,Y)$ for $\deg g=1$ satisfying $g(\alpha_{0},0)=0$ if and only if the linear system (2.2) has a solution in $k(\alpha_{0})$ for $m=1$.
@@
If (2.2) has a solution in $k(\alpha_{0})$ for $m=1$, then we have $\ell=d$ and there exists $g(X,Y)\in k(\alpha_{0})[X,Y]$, linear in $X$ and $Y$, such that $g(\alpha,Y)\equiv 0\pmod{Y^{\ell+1}}$. Consider $R(Y)=\mathrm{Res}_{X}(f,g)\in k(\alpha_{0})[Y]$, which is a polynomial in $Y$ of degree at most $d$. Since $f(\alpha,Y)\equiv 0\pmod{Y^{\ell+1}}$ as well, we have $R(Y)\equiv 0\pmod{Y^{\ell+1}}$, This forces $R=0$, so $f$ and $g$ have a nontrivial common factor in $k(\alpha_{0})[X,Y]$. But $g(X,Y)$ is already linear, so we have $g\mid f$ and $g(\alpha_{0},0)=0$.
Suppose $f(X,Y)=g(X,Y)h(X,Y)$ over $\overline{k}$ for $\deg g=1$ satisfying $g(\alpha_{0},0)=0$. Since $\alpha_{0}$ is a simple root of $f(X,0)$, we must in fact have $g$ defined over $k(\alpha_{0})$; otherwise, the conjugates of $g$ in $\overline{k}/k(\alpha_{0})$ would also vanish at $(\alpha_{0},0)$, producing a multiple root of $f(X,0)$ at $\alpha_{0}$. With $g,h$ defined over $k(\alpha_{0})$, we also have $h(\alpha_{0},0)\neq 0$, so $h(\alpha,Y)\not\equiv 0\pmod{Y}$. Since $f(\alpha,Y)\equiv 0\pmod{Y^{\ell_{\max}+1}}$, we have $g(\alpha,Y)\equiv 0\pmod{Y^{\ell_{\max}+1}}$. The coefficients of $g$ thus produce a solution to (2.2) in $k(\alpha_{0})$ for $m=1$. ∎
If $1<D\leq d-1$, we can still apply the strategy above to study the factorization of $f$. Given a solution to (2.2), we similarly produce $g$ of degree at most $D$ which has a common factor with $f$ and satisfies $g(\alpha_{0},0)=0$. However, when $D>1$, this does not guarantee that the common factor also vanishes on $(\alpha_{0},0)$.
For example, if $D=2$ and $f$ has a linear factor not vanishing on $(\alpha_{0},0)$, we could construct a solution to (2.2) with $m$ at most 2; we can no longer get away with degree 1, since we have asked for our factor $g$ to vanish on $(\alpha_{0},0)$. More generally for $D>1$, if (2.2) has no solutions in $\overline{k}$ for $m\leq D$, we may conclude that $f$ has no factors of degree at most $D$ vanishing on $(\alpha_{0},0)$, nor any factors of degree at most D1D-1. In particular, when D=d1D=d-1, this gives an irreducibility test which may be of independent interest; the proof is analogous to that of Proposition 2.4.
Proposition 2.5.
Then $f$ factors over $\overline{k}$ if and only if the linear system (2.2) has a solution in $k(\alpha_{0})$ for some $m<d$.
2.2. A genericity condition
Theorem 2.6.
Suppose $f(x_{1},\ldots,x_{n})$ is a (not necessarily homogeneous) polynomial of total degree $d$ over $k$, possessing no linear factors over $\overline{k}[x_{1},\ldots,x_{n}]$. Moreover, suppose there exist $(\mathbf{v}_{0},\mathbf{w}_{0})\in k^{2n-1}$ such that
$f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,0)=f(X+v_{1},w_{2}X+v_{2},\ldots,w_{n}X+v_{n})$
is a polynomial of degree $d$ in $k[X]$ with a simple root $\alpha_{0}\in\overline{k}$.
There exists $\Psi\in k(\alpha_{0})[Z_{2},\ldots,Z_{n}]$, depending on $\mathbf{v}_{0},\mathbf{w}_{0}$, of degree
$\deg\Psi\leq 3d-3$
such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,Y)$ has no linear factor over $\overline{k}$ vanishing on $(\alpha_{0},0)$ whenever $\Psi(z_{2},\ldots,z_{n})\neq 0$.
@@ 2.2. A genericity condition
We follow the same lines as [Kal95, Theorem 5] and [CM06, Theorem 3.3]. Let
$\chi=\chi(X,Y,Z_{2},\ldots,Z_{n})=f(X+v_{1},w_{2}X+Z_{2}Y+v_{2},\ldots,w_{n}X+Z_{n}Y+v_{n}),$
considered in $\overline{k}[X,Y,Z_{2},\ldots,Z_{n}]$. We first establish that $\chi$ has no factor linear in $X,Z_{2},\ldots,Z_{n}$, reminiscent of [Kal95, Lemma 5].
Suppose $\chi=\chi_{1}\chi_{2}$ with $\chi_{1}$ linear in $X,Z_{2},\ldots,Z_{n}$ over $\overline{k}[Y]$. Viewing $\chi\in\overline{k}[Y,Z_{2},\ldots,Z_{n}][X]$, the leading term in $X$ is a constant in $k$ (depending on $\mathbf{w}_{0}$) times $X^{d}$, by hypothesis. Thus both $\chi_{1}$ and $\chi_{2}$ must have positive degree in $X$.
Under the ring map
$\overline{k}[X,Y,Z_{2},\ldots,Z_{n}]$ $\to\overline{k}(Y)[x_{1},\ldots,x_{n}]$
$X$ $\mapsto x_{1}-v_{1}$
$Z_{i}$ $\mapsto\frac{1}{Y}\left(x_{i}-w_{i}(x_{1}-v_{1})-v_{i}\right)\text{ for }2\leq i\leq n,$
we see that $\chi$ maps to the image of $f$ under the inclusion $\overline{k}[x_{1},\ldots,x_{n}]\to\overline{k}(Y)[x_{1},\ldots,x_{n}]$. The map also preserves the total degree in $X,Z_{2},\ldots,Z_{n}$ (it is invertible upon inverting $Y$ in the domain), so the image of $\chi_{1}$ remains linear in $x_{1},\ldots,x_{n}$ over $\overline{k}(Y)$.
Thus we have that the image of $f$ in $\overline{k}(Y)[x_{1},\ldots,x_{n}]$ has a linear factor in the image of $\chi_{1}$. Since $f$ does not depend on $Y$ in the first place, this implies $f$ has a linear factor in $\overline{k}[x_{1},\ldots,x_{n}]$, contradicting the hypothesis.
Now consider $\chi\in\overline{k}(Z_{2},\ldots,Z_{n})[X,Y]$. Since we have established that $\chi$ has no linear factors, let alone any containing $(\alpha_{0},0)$, by Proposition 2.4 the system (2.2) has no solutions in $\overline{k}(Z_{2},\ldots,Z_{n})(\alpha_{0})$ for $m=1$. Moreover, upon examination of the Newton iteration step [Kal95, CM06, Step N] we observe the system (2.2) has coefficients in $k(\alpha_{0})[Z_{2},\ldots,Z_{n}]$; the only denominators introduced are those involving $\frac{\partial f}{\partial x}(\alpha_{0},0)\in k(\alpha_{0})$.
Let $\Psi$ be a nonzero determinant of a $3\times 3$ minor of the augmented matrix of (2.2). If $\mathbf{z}\in k^{n-1}$ satisfies $\Psi(z_{2},\ldots,z_{n})\neq 0$ then the system (2.2) obtained by substituting $Z_{i}=z_{i}$ for $2\leq i\leq n$ is inconsistent. Thus by Proposition 2.4, $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,Y)$ has no linear factors defined over $\overline{k}$ vanishing on $(\alpha_{0},0)$.
Following the degree computations for $\Psi_{1}^{(i)}$ in the proof of Theorem 2.1 (see also the proofs of [CM06, Theorem 3.3] and [Kal95, Theorem 5]), we find
$\deg\Psi\leq\sum_{j=0}^{2}\ell_{\max}-j=3d-3,$
completing the proof. ∎
Setting $k=\mathbb{F}_{q}$, and recognizing the number of $\mathbf{z}\in\mathbb{F}_{q}^{n-1}$ for which $\Psi(z_{2},\ldots,z_{n})=0$ is at most $(\deg\Psi)q^{n-2}$, we arrive at the following corollary.
Corollary 2.7.
Suppose $k=\mathbb{F}_{q}$ and $f$ satisfies the hypotheses of Theorem 2.6. Then there exist at most $(3d-3)q^{n-2}$ tuples $\mathbf{z}\in\mathbb{F}_{q}^{n-1}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}$ has no linear factor over $\overline{\mathbb{F}_{q}}$ containing $(\alpha_{0},0)$.
2.3. Proof of Theorem 2.1
We conclude this section by giving a proof of Theorem 2.1. We follow the approach of Cafure and Matera [CM06, §3] and highlight the minor modification to the necessary order of approximation which leads to the modest improvement over [CM06, Corollaries 3.2, 3.4]; see also [Kal95].

### v2

@@ 2.1. A factorization algorithm
In this subsection only, suppose $f(X,Y)\in k[X,Y]$ is monic in $X$ and has total degree $d$ (not necessarily homogeneous). Fix a positive integer $D<d$. The factorization algorithms found in [KAL95] and [CM06, §3.1] underlying the proofs of effective Bertini theorems [CM06, Corollaries 3.2, 3.3] request that $f(X,0)$ is squarefree. With slight modifications however, they still yield useful information when $f(X,0)$ is not squarefull, i.e. when considered over $\overline{k}$, $f(X,0)$ has at least one simple root. We recall the setup for the algorithm; see also [KAL95, §1], [CM06, §3.1].
@@
A solution to (2.2) for some $m\leq D$ implies that $f$ has a nonconstant factor of degree at most $D$. We establish this below, along with a partial converse. This is closely related to [CM06, Lemma 3.1], which puts stronger hypotheses on $f$ and uses a different order of approximation. [^1] In [CM06, §3.1] $\ell_{\max}=2dD$ and $\ell=2dm$ were used.
@@
Suppose $f(X,Y)\in k[X,Y]$ is a (not necessarily homogeneous) polynomial of total degree $d$ such that $f(X,0)$ is degree $d$ in $k[X]$ with a simple root $\alpha_{0}\in\overline{k}$. Fix $D<d$.
If the linear system (2.2) has a solution in $k(\alpha_{0})$ for $m\leq D$, then $f$ factors as $f=g(X,Y)h(X,Y)$ for nonconstant $g,h\in\overline{k}[X,Y]$ satisfying one of the following:
  1. (a)
    $\deg g\leq D$ and $g(\alpha_{0},0)=0$;
  2. (b)
    $\deg g\leq D-1$.
Conversely, if (a) above is satisfied, then the linear system (2.2) has a solution in $k(\alpha_{0})$ for some $m\leq D$.
Proof.
If (2.2) has a solution in $k(\alpha_{0})$ for $m\leq D$, then we have $\ell=dm$ and there exists $g(X,Y)\in k(\alpha_{0})[X,Y]$ of total degree at most $m$, such that $g(\alpha,Y)\equiv 0\pmod{Y^{\ell+1}}$. Consider $R(Y)=\mathrm{Res}_{X}(f,g)\in k(\alpha_{0})[Y]$, which is a polynomial in $Y$ of degree at most $\ell$. Since $f(\alpha,Y)\equiv 0\pmod{Y^{\ell+1}}$ as well, we have $R(Y)\equiv 0\pmod{Y^{\ell+1}}$. This forces $R=0$, so $f$ and $g$ have a nontrivial common factor in $k(\alpha_{0})[X,Y]$. Either $g\mid f$ and (a) is satisfied, or this common factor has degree strictly less than that of $g$, and hence less than $D$, so (b) is satisfied.
For the partial converse, suppose $f=gh$ over $\overline{k}$ satisfies (a). Since $\alpha_{0}$ is a simple root of $f(X,0)$, we have $g$ must be defined over $k(\alpha_{0})$; otherwise the conjugates of $g$ in $\overline{k}/k(\alpha_{0})$ would also vanish at $(\alpha_{0},0)$. Moreover, we have $h(\alpha_{0},0)\neq 0$, so $h(\alpha,Y)\not\equiv 0\pmod{Y^{\ell_{\max}+1}}$. Since $f(\alpha,Y)\equiv 0\pmod{Y^{\ell_{\max}+1}}$, we have $g(\alpha,Y)\equiv 0\pmod{Y^{\ell_{\max}+1}}$. The coefficients of $g$ thus produce a solution to (2.2) in $k(\alpha_{0})$ for $m=\deg g$. ∎
When $D=1$, Proposition 2.7 specializes to the statement that (2.2) has a solution in $k(\alpha_{0})$ for $m=1$ if and only if $f$ has a linear factor vanishing on $(\alpha_{0},0)$. We can also extract an equivalence in the case D=d1D=d-1.
Corollary 2.8.
@@
The linear system (2.2) has a solution in $k(\alpha_{0})$ for some $m<d$ if and only if $f=gh$ for nonconstant $g,h\in k(\alpha_{0})[X,Y]$.
2.2. Genericity conditions
Our goal is to constrain non-generic behavior of $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ to those $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$ that satisfy certain polynomial conditions. We reserve the lowercase $\beta,\gamma,\delta$ for elements of $k$, and use the uppercase $B,\Gamma,\Delta$ for indeterminates to describe these conditions. We give a flexible such result below; this is the key intermediate in proving Theorems 2.1 and 2.2.
Theorem 2.9.
Let $d\geq 2$ and $n\geq 3$. Suppose $f\in k[x_{0},\ldots,x_{n}]$ is a degree $d$ form, irreducible over $\overline{k}$. Moreover, suppose there exist $(\boldsymbol{\beta},\boldsymbol{\gamma})\in k^{2n-1}$ such that
$f(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})\in k[X]$
has degree $d$ with a simple root $\alpha_{0}\in\overline{k}$. Fix $0\leq D<d$.
There exists $\Psi_{D}^{(\alpha_{0})}\in k(\alpha_{0})[\Delta_{2},\ldots,\Delta_{n}]$ (depending on $(\boldsymbol{\beta},\boldsymbol{\gamma})$) of degree
$\deg\Psi_{D}^{(\alpha_{0})}\leq\frac{D}{8}\left(-D^{3}+4dD^{2}-6D^{2}+12dD-11D+8d-6\right)$
such that whenever $\Psi_{D}^{(\alpha_{0})}(\delta_{2},\ldots,\delta_{n})\neq 0$, $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ has no nonconstant factor of degree at most $D$ over $\overline{k}$ vanishing on $(\alpha_{0},0)$.
In particular, whenever $\Psi_{d-1}^{(\alpha_{0})}(\delta_{2},\ldots,\delta_{n})\neq 0$, $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ is irreducible over $\overline{k}$.
@@ 2.2. Genericity conditions
We follow similar lines as [KAL95, Theorem 5] and [CM06, Theorem 3.3]. Let
$\chi=\chi(X,Y,\Delta_{2},\ldots,\Delta_{n})=f(1,X+\beta_{1},\gamma_{2}X+\Delta_{2}Y+\beta_{2},\ldots,\gamma_{n}X+\Delta_{n}Y+\beta_{n}),$
considered in $\overline{k}[X,Y,\Delta_{2},\ldots,\Delta_{n}]$. We have $\chi$ is irreducible by [KAL95, Lemma 5].
Now consider $\chi\in\overline{k}(\Delta_{2},\ldots,\Delta_{n})[X,Y]$. By Proposition 2.7, the system (2.2) has no solutions in the field $\overline{k}(\Delta_{2},\ldots,\Delta_{n})$ for $m\leq D$. Moreover, upon examination of the Newton iteration step [KAL95, CM06, Step N] we observe the system (2.2) has coefficients in $k(\alpha_{0})[\Delta_{2},\ldots,\Delta_{n}]$; the only denominators introduced are those involving $\frac{\partial\chi}{\partial X}(\alpha_{0},0)\in k(\alpha_{0})$.
Let $\Psi_{D}^{(\alpha_{0})}$ be a maximal nonzero minor of the augmented matrix of (2.2). If $\boldsymbol{\delta}\in k^{n-1}$ is chosen such that $\Psi_{D}^{(\alpha_{0})}(\delta_{2},\ldots,\delta_{n})\neq 0$ then the system (2.2) obtained by specializing $\Delta_{i}$ to $\delta_{i}$ for $2\leq i\leq n$ is inconsistent. Thus by Proposition 2.7, $\chi(X,Y,\delta_{2},\ldots,\delta_{n})=f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ has no nonconstant factors of degree at most $D$ defined over $\overline{k}$ vanishing on $(\alpha_{0},0)$.
To estimate the degree of $\Psi_{D}^{(\alpha_{0})}$, the arguments of [CM06, Theorem 3.3] (see also [KAL95, Theorem 5]) remain essentially unchanged, save for the updated value of $\ell_{\max}=dD$. We find
$\deg\Psi_{D}^{(\alpha_{0})}\leq\sum_{j=0}^{(D+1)(D+2)/2-1}\ell_{\max}-j=\frac{D}{8}\left(-D^{3}+4dD^{2}-6D^{2}+12dD-11D+8d-6\right),$
completing the proof. ∎
We will also need that $f(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})$ is generically squarefree of degree $d$, and in fact this can be controlled by another polynomial condition. This is [KAL95, Lemma 4], modified to allow reducible $f$.
Lemma 2.10.
Let $d\geq 2$ and $n\geq 3$. Suppose $f\in k[x_{0},\ldots,x_{n}]$ is a degree $d$ form with factorization type $(d_{1},\ldots,d_{t})$ over $\overline{k}$ such that $x_{0}\nmid f$.
There exists a nonzero polynomial $\Upsilon\in\overline{k}[B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n}]$ with $\deg\Upsilon\leq 2d^{2}$ such that whenever $\Upsilon(\beta_{1},\ldots,\beta_{n},\gamma_{2},\ldots,\gamma_{n})\neq 0$ for $(\boldsymbol{\beta},\boldsymbol{\gamma})\in\overline{k}^{2n-1}$, we have $f(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})\in\overline{k}[X]$ is squarefree of degree $d$.
@@
We essentially follow [CM06, §3.2] and [KAL95, Lemmas 2,3,4]. Let
$\phi=f(1,X+B_{1},\Gamma_{2}X+Y_{2}+B_{2},\ldots,\Gamma_{n}X+Y_{n}+B_{n})\in k(B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n})[X,Y_{2},\ldots,Y_{n}]$
and similarly define $\phi_{i}$ in terms of the irreducible factors $f_{i}$. By [KAL95, Lemma 2], the $\phi_{i}$ are irreducible over the algebraic closure of $k(B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n})$. The same argument in [KAL95, Lemma 2] shows $\phi_{i}$ and $\phi_{j}$ share no common factors over the closure of $k(B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n})$, since otherwise so would the $f_{i}$.
Let $\lambda$ (resp. $\lambda_{i}$) denote the leading term of $\phi(X,0,\ldots,0)$ (resp. $\phi_{i}(X,0,\ldots,0)$) as a polynomial in $k(B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n})[X]$. We have $\lambda,\lambda_{i}\in k[\Gamma_{2},\ldots,\Gamma_{n}]$ have total degree $d$ and $d_{i}$, respectively (the hypothesis $x_{0}\nmid f$ ensures that $\phi(X,0,\ldots,0)$ has degree $d$ in $X$).
Take $\psi=\phi/\lambda$ and $\psi_{i}=\phi_{i}/\lambda_{i}$, which are in $k(B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n})[X,Y_{2},\ldots,Y_{n}]$, and now monic in $X$. We want to show the resultant $\operatorname{Res}_{X}\left(\psi(X,0,\ldots,0),\frac{\partial\psi(X,0,\ldots,0)}{\partial X}\right)$ is nonzero in $k(B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n})$, which generically gives the desired squarefreeness. By basic properties of resultants,
$\operatorname{Res}_{X}$ $\Big(\psi(X,0,\ldots,0),\frac{\partial\psi(X,0,\ldots,0)}{\partial X}\Big)$
$=\prod_{1\leq i\leq t}\operatorname{Res}_{X}\left(\psi_{i}(X,0,\ldots,0),\frac{\partial\psi_{i}(X,0,\ldots,0)}{\partial X}\right)\prod_{j\neq i}\operatorname{Res}_{X}\left(\psi_{i}(X,0,\ldots,0),\psi_{j}(X,0,\ldots,0)\right).$
By [KAL95, Lemma 3], we have that $\operatorname{Res}_{X}\left(\psi_{i}(X,0,\ldots,0),\frac{\partial\psi_{i}(X,0,\ldots,0)}{\partial X}\right)$ is nonzero. Furthermore, following the strategy of the proof of [KAL95, Lemma 3], we find that if $\operatorname{Res}_{X}\left(\psi_{i}(X,0,\ldots,0),\psi_{j}(X,0,\ldots,0)\right)\neq 0$, then there is a common nonconstant factor $g(X)\in k(\Gamma_{2},\ldots,\Gamma_{n})[X,B_{1},\ldots,B_{n}]$ of $\psi_{i}(X,0,\ldots,0)$ and $\psi_{j}(X,0,\ldots,0)$. Making the substitution
$g\left(x_{1}-B_{1},B_{1},x_{2}-\Gamma_{2}(x_{1}-B_{1}),\ldots,x_{n}-\Gamma_{n}(x_{1}-B_{1})\right),$
and rehomogenizing, we obtain a nonconstant factor of both $f_{i}$ and $f_{j}$ over $k$, a contradiction. Thus we conclude the desired resultant is nonzero, from which it follows that $f(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})$ (and hence $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,0)$) is generically squarefree for $(\boldsymbol{\beta},\boldsymbol{\gamma})\in\overline{k}^{2n-1}$.
Take $\Upsilon$ to be the product of $\lambda$ and $\operatorname{Res}_{X}\left(\phi(X,0,\ldots,0),\frac{\partial\phi(X,0,\ldots,0)}{\partial X}\right)$. When $\Upsilon(\beta_{1},\ldots,\beta_{n},\gamma_{2},\ldots,\gamma_{n})\neq 0$, we have that
$f(1,X+\beta_{1},\gamma_{2}X+\beta_{2},\ldots,\gamma_{n}X+\beta_{n})$
has nonzero leading term and discriminant, i.e. it is squarefree of degree $d$ in $\overline{k}[X]$.
It remains to compute the degree of $\operatorname{Res}_{X}\left(\phi(X,0,\ldots,0),\frac{\partial\phi(X,0,\ldots,0)}{\partial X}\right)$ in $k[B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n}]$. As in [KAL95, Lemma 4], this follows from the fact that the resultant is homogeneous of degree 2d12d-1 in the coefficients of $\phi(X,0,\ldots,0)$ and $\frac{\partial\phi(X,0,\ldots,0)}{\partial X}$; these coefficients have total degree at most $d$ in $B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n}$, making the total degree of the generic resultant at most $2d^{2}-d$. Adding in the degree of $\lambda$, we have $\deg\Upsilon\leq 2d^{2}$. ∎
2.3. Proofs of Theorem 2.1 and 2.2